# The algebraic and geometric classification of

### v1

@@
Abstract: In this paper, we develop a method to obtain the algebraic classification of compatible pre-Lie algebras from the classification of pre-Lie algebras of the same dimension. We use this method to obtain the algebraic classification of complex $2$-dimensional compatible pre-Lie algebras. As a byproduct, we obtain the classification of complex $2$-dimensional compatible commutative associative, compatible associative and compatible Novikov algebras. In addition, we consider the geometric classification of varieties of cited algebras, that is the description of its irreducible components.
Keywords: compatible algebra, compatible associative algebra, compatible pre-Lie algebra, algebraic classification, geometric classification.
MSC2020: 17A30 (primary); 17D25, 14L30 (secondary).
Introduction
@@
The algebraic classification (up to isomorphism) of algebras of small dimensions from a certain variety defined by a family of polynomial identities is a classic problem in the theory of non-associative algebras. Another interesting approach to studying algebras of a fixed dimension is to study them from a geometric point of view (that is, to study the degenerations and deformations of these algebras). The results in which the complete information about degenerations of a certain variety is obtained are generally referred to as the geometric classification of the algebras of these varieties. There are many results related to the algebraic and geometric classification of Jordan, Lie, Leibniz, Zinbiel, and other algebras (see, [7, 6, 28, 1, 4, 5, 18, 3, 25, 27, 16, 17, 23, 2] and references in [24, 32, 26]).
@@
Pre-Lie algebras, also known as right symmetric algebras, appeared in some papers by Gerstenhaber, Koszul, and Vinberg in the 1960s. It is a generalization of associative algebras and the most popular non-associative subvariety of Lie-admissible algebras. They have various applications in geometry and physics [8]; recently some connections between them and trees, braces and $F$-manifold algebras were established in [38, 37, 9, 30]. Novikov algebras are pre-Lie algebras with an additional identity. They were introduced in papers by Gel’fand and Dorfman, Balinskiĭ and Novikov (about Novikov algebras see [15] and references therein). Let $\Omega$ be a variety of algebras. We say that an algebra $(\bf A,\cdot,\ast)$ is a compatible $\Omega$-algebra, if and only if $(\bf A,\cdot),$ $(\bf A,\ast)$ and $(\bf A,\cdot+\ast)$ are $\Omega$-algebras. Compatible Lie algebras are considered in the study of the classical Yang-Baxter equation [21], integrable equations of the principal chiral model type [20], elliptic theta functions [33], and other areas. The study of non-Lie compatible algebras is also very popular. So, cohomology and deformations were studied for compatible Lie algebras [31], compatible associative algebras [10], compatible ${\rm Hom}$-Lie algebras [12], compatible $L_{\infty}$-algebras [11], compatible $3$-Lie algebras [19], compatible dendriform algebras [13], and so on. Free compatible algebras were studied in the associative algebra case in [14] and in the Lie algebra case in [22, 29]. Compatible associative structures on matrix algebras were studied in [34, 35, 36]. General compatible structures have been studied from an operadic point of view in [39, 40]. At this moment there is only one paper about the classification of small dimensional compatible algebras. Namely, all $4$-dimensional nilpotent compatible Lie algebras were classified in [28].
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible pre-Lie algebras, described explicitly in Theorem 7 in terms of $6$ three-parameter families, $14$ two-parameter families, $13$ one-parameter families, and $8$ additional isomorphism classes.
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible commutative associative algebras, described explicitly in Theorem 8 in terms of $1$ three-parameter family, $3$ two-parameter families, $8$ one-parameter families, and $6$ additional isomorphism classes.
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible associative algebras, described explicitly in Theorem 10 in terms of $1$ three-parameter family, $3$ two-parameter families, $10$ one-parameter families, and $10$ additional isomorphism classes.
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible Novikov algebras, described explicitly in Theorem 12 in terms of $3$ three-parameter family, $7$ two-parameter families, $11$ one-parameter families, and $7$ additional isomorphism classes.
@@
The variety of complex $2$-dimensional compatible pre-Lie algebras has dimension $7$. It is defined by $2$ rigid algebras, $1$ one-parametric family of algebras, $7$ two-parametric family of algebras, and $4$ three-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 22.
@@
The variety of complex $2$-dimensional compatible commutative associative algebras has dimension $7$. It is defined by $1$ two-parametric family of algebras and $1$ three-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 19.
@@
The variety of complex $2$-dimensional compatible associative algebras has dimension $7$. It is defined by $2$ rigid algebras, $2$ two-parametric family of algebras, and $1$ three-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 20.
@@
The variety of complex $2$-dimensional compatible Novikov algebras has dimension $7$. It is defined by $1$ rigid algebra, $3$ two-parametric family of algebras, and $1$ three-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 21.
1. The algebraic classification of compatible algebras
@@
1.1. The algebraic classification of algebras
@@
$\left(x\cdot y\right)\cdot z-x\cdot\left(y\cdot z\right)$ $=$ $\left(y\cdot x\right)\cdot z-y\cdot\left(x\cdot z\right).$
Definition 2.
A compatible pre-Lie algebra is a vector space ${\bf A}$ equipped with two multiplication $\cdot$ and another multiplication $\ast$,∗ , such that, $({\bf A},\cdot),$ $({\bf A},\ast)$ and $({\bf A},\cdot+\ast)$ are pre-Lie algebras. These two operations are required to satisfy the following identities.
$\left(x\cdot y\right)\cdot z-x\cdot\left(y\cdot z\right)$ $=$ $\left(y\cdot x\right)\cdot z-y\cdot\left(x\cdot z\right),$
$\left(x\ast y\right)\ast z-x\ast\left(y\ast z\right)$ $=$ $\left(y\ast x\right)\ast z-y\ast\left(x\ast z\right),$
$(x\ast y)\cdot z-x\ast(y\cdot z)+(x\cdot y)\ast z-x\cdot(y\ast z)$ $=$ $(y\ast x)\cdot z-y\ast(x\cdot z)+(y\cdot x)\ast z-y\cdot(x\ast z).$
@@
Definition 3.
Let $\left({\bf A},\cdot\right)$ be a pre-Lie algebra. Define ${\rm Z}^{2}\left({\bf A},{\bf A}\right)$ to be the set of all bilinear maps $\theta:{\bf A}\times{\bf A}\longrightarrow{\bf A}$ such that:
$\theta(\theta\left(x,y\right),z)-\theta(x,\theta\left(y,z\right))$ $=$ $\theta(\theta\left(y,x\right),z)-\theta(y,\theta\left(x,z\right)),$
$\theta(x,y)\cdot z-\theta(x,y\cdot z)+\theta(x\cdot y,z)-x\cdot\theta(y,z)$ $=$ $\theta(y,x)\cdot z-\theta(y,x\cdot z)+\theta(y\cdot x,z)-y\cdot\theta(x,z).$
Then ${\rm Z}^{2}\left({\bf A},{\bf A}\right)\neq{\varnothing}$ since $\theta=0\in{\rm Z}^{2}\left({\bf A},{\bf A}\right)$.
Now, for $\theta\in{\rm Z}^{2}\left({\bf A},{\bf A}\right)$, let us define a multiplication $\bullet_{\theta}$ on ${\bf A}$ by $x\bullet_{\theta}y=\theta\left(x,y\right)$ for all $x,y\in{\bf A}$. Then $\left({\bf A},\cdot,\bullet_{\theta}\right)$ is a compatible pre-Lie algebra. Conversely, if $\left({\bf A},\cdot,\bullet\right)$ is a compatible pre-Lie algebra, then there exists $\theta\in{\rm Z}^{2}\left({\bf A},{\bf A}\right)$ such that $\left({\bf A},\cdot,\bullet_{\theta}\right)\cong\left({\bf A},\cdot,\bullet\right)$. To see this, consider the bilinear map $\theta:{\bf A}\times{\bf A}\longrightarrow{\bf A}$ defined by $\theta\left(x,y\right)=x\bullet y$ for all $x,y$ in ${\bf A}$. Then $\theta\in{\rm Z}^{2}\left({\bf A},{\bf A}\right)$ and $\left({\bf A},\cdot,\bullet_{\theta}\right)=\left({\bf A},\cdot,\bullet\right)$.
Let $\left({\bf A},\cdot\right)$ be a pre-Lie algebra. The automorphism group, ${\rm Aut}\left({\bf A}\right)$, of ${\bf A}$ acts on ${\rm Z}^{2}\left({\bf A},{\bf A}\right)$ by
$\left(\theta*\phi\right)\left(x,y\right)$ $=$ $\phi^{-1}\left(\theta\left(\phi\left(x\right),\phi\left(y\right)\right)\right),$
for $\phi\in{\rm Aut}\left({\bf A}\right)$, and $\theta\in{\rm Z}^{2}\left({\bf A},{\bf A}\right)$.
Lemma 4.
Let $\left({\bf A},\cdot\right)$ be a pre-Lie algebra and $\theta,\vartheta\in{\rm Z}^{2}\left({\bf A},{\bf A}\right)$. Then $\left({\bf A},\cdot,\bullet_{\theta}\right)$ and $\left({\bf A},\cdot,\bullet_{\vartheta}\right)$ are isomorphic if and only if there is a $\phi\in{\rm Aut}\left({\bf A}\right)$ with $\theta*\phi=\vartheta$.
@@
If $\theta*\phi=\vartheta$, then $\phi:\left({\bf A},\cdot,\bullet_{\vartheta}\right)\longrightarrow$ $\left({\bf A},\cdot,\bullet_{\theta}\right)$ is an isomorphism since $\phi\left(\vartheta\left(x,y\right)\right)=\theta\left(\phi\left(x\right),\phi\left(y\right)\right)$. On the other hand, if $\phi:\left({\bf A},\cdot,\bullet_{\vartheta}\right)\longrightarrow$ $\left({\bf A},\cdot,\bullet_{\theta}\right)$ is an isomorphism of compatible pre-Lie algebras, then $\phi\in{\rm Aut}\left({\bf A}\right)$ and $\phi\left(x\bullet_{\vartheta}y\right)=\phi\left(x\right)\bullet_{\theta}\phi\left(y\right)$. Hence
$\vartheta\left(x,y\right)$ $=$ $\phi^{-1}\left(\theta\left(\phi\left(x\right),\phi\left(y\right)\right)\right)$ $=$ $\left(\theta*\phi\right)\left(x,y\right)$
@@
Consequently, we have a procedure to classify the compatible pre-Lie algebras with given associated pre-Lie algebra $\left({\bf A},\cdot\right)$. It consists of three steps:
@@
    Find the orbits of ${\rm Aut}\left({\bf A}\right)$ on ${\rm Z}^{2}\left({\bf A},{\bf A}\right)$.
@@
    Choose a representative $\theta$ from each orbit and then construct the compatible pre-Lie algebra $\left({\bf A},\cdot,\bullet_{\theta}\right)$.
1.2. $2$-dimensional pre-Lie algebras
Lemma 5.
@@ 1.2. $2$-dimensional pre-Lie algebras
$\mathcal{C}_{01}$ :::: $e_{1}\cdot e_{1}=e_{1}+e_{2},$ $e_{2}\cdot e_{1}=e_{2}.$
$\mathcal{C}_{02}$ :::: $e_{1}\cdot e_{1}=e_{1}+e_{2},$ $e_{1}\cdot e_{2}=e_{2}.$
$\mathcal{C}_{03}$ :::: $e_{1}\cdot e_{1}=e_{2}.$
$\mathcal{C}_{04}$ :::: $e_{2}\cdot e_{1}=e_{1}.$
$\mathcal{C}_{05}^{\alpha}$ :::: $e_{1}\cdot e_{1}=e_{1},$ $e_{1}\cdot e_{2}=\alpha e_{2}.$
$\mathcal{C}_{06}^{\alpha}$ :::: $e_{1}\cdot e_{1}=e_{1},$ $e_{1}\cdot e_{2}=\alpha e_{2},$ $e_{2}\cdot e_{1}=e_{2}.$
$\mathcal{C}_{07}$ :::: $e_{1}\cdot e_{1}=e_{1},$ $e_{2}\cdot e_{2}=e_{2}.$
$\mathcal{C}_{08}$ :::: $e_{1}\cdot e_{1}=e_{1},$ $e_{1}\cdot e_{2}=2e_{2},$ $e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2},$ $e_{2}\cdot e_{2}=e_{2}.$
Lemma 6.
The description of the group of automorphisms of every $2$-dimensional pre-Lie algebra is given below.

### v2

@@
[…]
The Algebraic and Geometric Classification of Compatible Pre-Lie Algebras
[…]
The Algebraic and Geometric Classification
of Compatible Pre-Lie Algebras
[…]
Hani ABDELWAHAB a, Ivan KAYGORODOV b and Abdenacer MAKHLOUF c
[…]
H. Abdelwahab, I. Kaygorodov and A. Makhlouf
[…]
Received August 22, 2024, in final form November 18, 2024; Published online November 28, 2024
[…]
@@
In this paper, we develop a method to obtain the algebraic classification of compatible pre-Lie algebras from the classification of pre-Lie algebras of the same dimension. We use this method to obtain the algebraic classification of complex 2-dimensional compatible pre-Lie algebras. As a byproduct, we obtain the classification of complex 2-dimensional compatible commutative associative, compatible associative and compatible Novikov algebras. In addition, we consider the geometric classification of varieties of cited algebras, that is the description of its irreducible components.
[…]
compatible algebra; compatible associative algebra; compatible pre-Lie algebra; algebraic classification; geometric classification
[…]
17A30; 17D25; 14L30
1 Introduction
@@
The algebraic classification (up to isomorphism) of algebras of small dimensions from a certain variety defined by a family of polynomial identities is a classic problem in the theory of non-associative algebras. Another interesting approach to studying algebras of a fixed dimension is to study them from a geometric point of view (that is, to study the degenerations and deformations of these algebras). The results in which the complete information about degenerations of a certain variety is obtained are generally referred to as the geometric classification of the algebras of these varieties. There are many results related to the algebraic and geometric classification of Jordan, Lie, Leibniz, Zinbiel, and other algebras (see [1, 2, 3, 5, 6, 15, 16, 20, 23, 25, 27] and references in [22, 24, 31]). The geometric classification of algebras from a certain variety is based on the notion of degeneration, that is a “dual” notion to deformation [4, 11, 15, 21, 29, 32].
@@
Pre-Lie algebras, also known as right symmetric algebras, appeared in some papers by Gerstenhaber, Koszul, and Vinberg in the 1960s. It is a generalization of associative algebras and the most popular non-associative subvariety of Lie-admissible algebras. They have various applications in geometry and physics [7]; recently some connections between them and trees, braces and $F$-manifold algebras were established in [8, 29, 37, 38]. Novikov algebras are pre-Lie algebras with an additional identity. They were introduced in papers by Gel’fand and Dorfman, Balinskii and Novikov (about Novikov algebras, see [14] and references therein). Let $\Omega$ be a variety of algebras. We say that an algebra $(\bf A,\cdot,\ast)$ is a compatible $\Omega$-algebra, if and only if $(\bf A,\cdot)$, $(\bf A,\ast)$ and $(\bf A,\cdot+\ast)$ are $\Omega$-algebras. Compatible Lie algebras are considered in the study of the classical Yang–Baxter equation [18], integrable equations of the principal chiral model type [17], elliptic theta functions [34], and other areas. The study of non-Lie compatible algebras is also very popular. So, cohomology and deformations were studied for compatible Lie algebras [30], compatible associative algebras [9], compatible ${\rm Hom}$-Lie algebras [11], compatible $L_{\infty}$-algebras [10], compatible $3$-Lie algebras [21], compatible dendriform algebras [12], and so on. Free compatible algebras were studied in the associative algebra case in [13] and in the Lie algebra case in [19, 28]. Compatible associative structures on matrix algebras were studied in [33, 35, 36]. General compatible structures have been studied from an operadic point of view in [39, 40]. A generalization of compatible algebras was considered in [26]. At this moment there is only one paper about the classification of small-dimensional compatible algebras. Namely, all $4$-dimensional nilpotent compatible Lie algebras were classified in [27].
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible pre-Lie algebras, described explicitly in Theorem 2.7 in terms of $6$ three-parameter families, $14$ two-parameter families, $13$ one-parameter families, and $8$ additional isomorphism classes.
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible commutative associative algebras, described explicitly in Theorem 2.8 in terms of $1$ three-parameter family, $3$ two-parameter families, $8$ one-parameter families, and $6$ additional isomorphism classes.
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible associative algebras, described explicitly in Theorem 2.10 in terms of $1$ three-parameter family, $3$ two-parameter families, $10$ one-parameter families, and $10$ additional isomorphism classes.
@@
There are infinitely many isomorphism classes of complex $2$-dimensional compatible Novikov algebras, described explicitly in Theorem 2.12 in terms of $3$ three-parameter families, $7$ two-parameter families, $11$ one-parameter families, and $7$ additional isomorphism classes.
@@
The variety of complex $2$-dimensional compatible pre-Lie algebras has dimension $7$. It is defined by $2$ rigid algebras, $1$ one-parametric family of algebras, $7$ two-parametric families of algebras, and $4$ three-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 3.10.
@@
The variety of complex $2$-dimensional compatible commutative associative algebras has dimension $7$. It is defined by $1$ two-parametric family of algebras and $1$ three-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 3.7.
@@
The variety of complex $2$-dimensional compatible associative algebras has dimension $7$. It is defined by $2$ rigid algebras, $2$ two-parametric families of algebras, and $1$ three-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 3.8.
@@
The variety of complex $2$-dimensional compatible Novikov algebras has dimension $7$. It is defined by $1$ rigid algebra, $3$ two-parametric families of algebras, and $1$ three-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{2}(\mathbb{C})$-orbits of the algebras given in Theorem 3.9.
2 The algebraic classification of compatible algebras
@@
2.1 The algebraic classification of algebras
@@
$(x\cdot y)\cdot z-x\cdot(y\cdot z)=(y\cdot x)\cdot z-y\cdot(x\cdot z).$
Definition 2.2.
A compatible pre-Lie algebra is a vector space ${\bf A}$ equipped with two multiplications: $\cdot$ and another multiplication $\ast$ such that $({\bf A},\cdot)$, $({\bf A},\ast)$ and $({\bf A},\cdot+\ast)$ are pre-Lie algebras. These two operations are required to satisfy the following identities:
$(x\cdot y)\cdot z-x\cdot(y\cdot z)=(y\cdot x)\cdot z-y\cdot(x\cdot z),$
$(x\ast y)\ast z-x\ast(y\ast z)=(y\ast x)\ast z-y\ast(x\ast z),$
$(x\ast y)\cdot z-x\ast(y\cdot z)+(x\cdot y)\ast z-x\cdot(y\ast z)$
$\qquad=(y\ast x)\cdot z-y\ast(x\cdot z)+(y\cdot x)\ast z-y\cdot(x\ast z).$
@@
Definition 2.3.
Let $({\bf A},\cdot)$ be a pre-Lie algebra. Define ${\rm Z}^{2}({\bf A},{\bf A})$ to be the set of all bilinear maps $\theta\colon{\bf A}\times{\bf A}\longrightarrow{\bf A}$ such that
$\theta(\theta(x,y),z)-\theta(x,\theta(y,z))=\theta(\theta(y,x),z)-\theta(y,\theta(x,z)),$
$\theta(x,y)\cdot z-\theta(x,y\cdot z)+\theta(x\cdot y,z)-x\cdot\theta(y,z)$
$\qquad=\theta(y,x)\cdot z-\theta(y,x\cdot z)+\theta(y\cdot x,z)-y\cdot\theta(x,z).$
Then ${\rm Z}^{2}({\bf A},{\bf A})\neq{\varnothing}$ since $\theta=0\in{\rm Z}^{2}({\bf A},{\bf A})$.
Now, for $\theta\in{\rm Z}^{2}({\bf A},{\bf A})$, let us define a multiplication $\bullet_{\theta}$ on ${\bf A}$ by $x\bullet_{\theta}y=\theta(x,y)$ for all ${x,y\in{\bf A}}$. Then $({\bf A},\cdot,\bullet_{\theta})$ is a compatible pre-Lie algebra. Conversely, if $({\bf A},\cdot,\bullet)$ is a compatible pre-Lie algebra, then there exists $\theta\in{\rm Z}^{2}({\bf A},{\bf A})$ such that $({\bf A},\cdot,\bullet_{\theta})\cong({\bf A},\cdot,\bullet)$. To see this, consider the bilinear map $\theta\colon{\bf A}\times{\bf A}\longrightarrow{\bf A}$ defined by $\theta(x,y)=x\bullet y$ for all $x,y$ in ${\bf A}$. Then $\theta\in{\rm Z}^{2}({\bf A},{\bf A})$ and $({\bf A},\cdot,\bullet_{\theta})=({\bf A},\cdot,\bullet)$.
Let $({\bf A},\cdot)$ be a pre-Lie algebra. The automorphism group ${\rm Aut}({\bf A})$ of ${\bf A}$ acts on ${\rm Z}^{2}({\bf A},{\bf A})$ by
$(\theta*\phi)(x,y)=\phi^{-1}(\theta(\phi(x),\phi(y)))$
for $\phi\in{\rm Aut}({\bf A})$ and $\theta\in{\rm Z}^{2}({\bf A},{\bf A})$.
Lemma 2.4.
Let $({\bf A},\cdot)$ be a pre-Lie algebra and $\theta,\vartheta\in{\rm Z}^{2}({\bf A},{\bf A})$. Then $({\bf A},\cdot,\bullet_{\theta})$ and $({\bf A},\cdot,\bullet_{\vartheta})$ are isomorphic if and only if there is a linear map $\phi\in{\rm Aut}({\bf A})$ such that $\theta*\phi=\vartheta$.
@@
If $\theta*\phi=\vartheta$, then $\phi\colon({\bf A},\cdot,\bullet_{\vartheta})\longrightarrow$ $({\bf A},\cdot,\bullet_{\theta})$ is an isomorphism since $\phi(\vartheta(x,y))=\theta(\phi(x),\phi(y))$. On the other hand, if $\phi\colon({\bf A},\cdot,\bullet_{\vartheta})\longrightarrow$ $({\bf A},\cdot,\bullet_{\theta})$ is an isomorphism of compatible pre-Lie algebras, then $\phi\in{\rm Aut}({\bf A})$ and $\phi(x\bullet_{\vartheta}y)=\phi(x)\bullet_{\theta}\phi(y)$. Hence
$\vartheta(x,y)=\phi^{-1}(\theta(\phi(x),\phi(y)))=(\theta*\phi)(x,y),$
@@
Consequently, we have a procedure to classify the compatible pre-Lie algebras with the given associated pre-Lie algebra $({\bf A},\cdot)$. It consists of three steps:
@@
    find the orbits of ${\rm Aut}({\bf A})$ on ${\rm Z}^{2}({\bf A},{\bf A})$,
@@
    choose a representative $\theta$ from each orbit and then construct the compatible pre-Lie algebra $({\bf A},\cdot,\bullet_{\theta})$.
2.2 2-dimensional pre-Lie algebras
Lemma 2.5.
@@
$\mathcal{C}_{01}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\mathcal{C}_{02}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{1}\cdot e_{2}=e_{2},$
$\mathcal{C}_{03}\colon\ e_{1}\cdot e_{1}=e_{2},$
$\mathcal{C}_{04}\colon\ e_{2}\cdot e_{1}=e_{1},$
$\mathcal{C}_{05}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},$
$\mathcal{C}_{06}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\mathcal{C}_{07}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},$
$\mathcal{C}_{08}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=2e_{2},\quad e_{2}\cdot e_{1}=\tfrac{1}{2}e_{1}+e_{2},\quad e_{2}\cdot e_{2}=e_{2}.$
Lemma 2.6.
The description of the group of automorphisms of every $2$-dimensional pre-Lie algebra is given as follows: